# Turb-L1: Achieving Long-term Turbulence Tracing By Tackling Spectral Bias

### v2

@@ 3.2 Main results
Table 1 presents the performance achieved across different baselines over various prediction steps. We summarize the key takeaways as follows:
[…]
Takeaway ➋: Turb-L1 demonstrates excellent visual fidelity and physical consistency in long-term predictions. It effectively overcomes over-smoothing and simulation collapse, issues that spectral bias causes. As Figure 3 illustrates, Turb-L1’s predictions closely match the ground truth even after 99 time steps, clearly reproducing complex vortex structures and fine flow features. This performance stems from its core Hierarchical Dynamics Synthesis mechanism, which precisely captures full-spectrum dynamic information, particularly high-frequency details. In contrast, FNO exhibits significant over-smoothing due to spectral bias and loses critical high-frequency information. SimVP, on the other hand, experiences physics collapse from accumulated high-frequency error amplification, producing unrealistic artifacts. Turb-L1’s SSIM and relative L2 error curves further confirm its significant advantages in maintaining structural similarity and long-term prediction accuracy. This fully demonstrates the method’s robustness and high fidelity in tracking complex multi-scale turbulence dynamics. Further ablation experiments are presented in the Appendix.
@@ 3.4 Turb-L1’s Superiority in Enstrophy Spectrum Prediction
4 Conclusion
Supplementary materials
[…]
1
Appendix A Proof of Theorem on High-Frequency Error Control
This section provides a detailed proof for Theorem A.1, which establishes a bound on the accumulation of high-frequency errors in long-term autoregressive predictions.
A.1 Preliminaries: Error Evolution Equation
Let the true state evolution of a dynamical system be governed by:
$\bm{\omega}(t_{k+1})=S(\bm{\omega}(t_{k})),$ (A.14)
where $\bm{\omega}(t_{k})$ is the true state at discrete time step $t_{k}$, and $S$ is the true evolution operator. A surrogate model approximates this evolution as:
$\hat{\bm{\omega}}(t_{k+1})=\mathcal{F}(\hat{\bm{\omega}}(t_{k})),$ (A.15)
where $\hat{\bm{\omega}}(t_{k})$ is the model-predicted state, and $\mathcal{F}$ is the learned model operator.
The prediction error at time $t_{k}$ is defined as $\bm{\epsilon}(t_{k})=\hat{\bm{\omega}}(t_{k})-\bm{\omega}^{*}(t_{k})$, where $\bm{\omega}^{*}(t_{k})$ denotes the true trajectory. The error at the next time step $t_{k+1}$ is:
$\bm{\epsilon}(t_{k+1})$ $=\hat{\bm{\omega}}(t_{k+1})-\bm{\omega}^{*}(t_{k+1})$ (A.16)
$=\mathcal{F}(\hat{\bm{\omega}}(t_{k}))-S(\bm{\omega}^{*}(t_{k})).$ (A.17)
Performing a first-order Taylor expansion of $\mathcal{F}(\hat{\bm{\omega}}(t_{k}))$ around the true state $\bm{\omega}^{*}(t_{k})$, assuming $\bm{\epsilon}(t_{k})$ is sufficiently small:
$\mathcal{F}(\hat{\bm{\omega}}(t_{k}))=\mathcal{F}(\bm{\omega}^{*}(t_{k})+\bm{\epsilon}(t_{k}))\approx\mathcal{F}(\bm{\omega}^{*}(t_{k}))+\nabla\mathcal{F}|_{\bm{\omega}^{*}(t_{k})}\bm{\epsilon}(t_{k}).$ (A.18)
Substituting this back into the error equation:
$\bm{\epsilon}(t_{k+1})\approx\mathcal{F}(\bm{\omega}^{*}(t_{k}))-S(\bm{\omega}^{*}(t_{k}))+\nabla\mathcal{F}|_{\bm{\omega}^{*}(t_{k})}\bm{\epsilon}(t_{k}).$ (A.19)
Let $\bm{e}_{\text{model}}(t_{k})=\mathcal{F}(\bm{\omega}^{*}(t_{k}))-S(\bm{\omega}^{*}(t_{k}))$ be the single-step model error. Assuming the Jacobian of the learned model $\nabla\mathcal{F}|_{\bm{\omega}^{*}(t_{k})}$ approximates the Jacobian of the true system $\mathbf{A}_{k}=\nabla S|_{\bm{\omega}^{*}(t_{k})}$, the error evolution is:
$\bm{\epsilon}(t_{k+1})\approx\mathbf{A}_{k}\bm{\epsilon}(t_{k})+\bm{e}_{\text{model}}(t_{k}).$ (A.20)
A.2 Theorem and Proof
Theorem A.1 (High-Frequency Error Control and Prediction Stability).
Assume the system Jacobian $\mathbf{A}_{k}=\nabla S|_{\bm{\omega}^{*}(t_{k})}$ amplifies high-frequency errors, with constants $C_{A}\geq 1$ and $\gamma_{\text{high}}>1$, such that for the high-frequency component of the error $\bm{\epsilon}_{\text{high}}(t_{k})=\mathbf{P}_{\text{high}}\bm{\epsilon}(t_{k})$ (where $\mathbf{P}_{\text{high}}$ is the projection operator onto a high-frequency subspace):
$\|\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}_{\text{high}}(t_{k})\|\leq C_{A}\gamma_{\text{high}}\|\bm{\epsilon}_{\text{high}}(t_{k})\|.$ (A.21)
$\|\bm{\epsilon}_{\text{high}}(t_{M})\|\leq(C_{A}\gamma_{\text{high}})^{M}\|\bm{\epsilon}_{\text{high}}(t_{0})\|+\delta_{\text{HF}}\frac{(C_{A}\gamma_{\text{high}})^{M}-1}{C_{A}\gamma_{\text{high}}-1},$ (A.22)
for $C_{A}\gamma_{\text{high}}\neq 1$ and initial high-frequency error $\bm{\epsilon}_{\text{high}}(t_{0})$.
Proof.
Let $\bm{\epsilon}_{\text{high}}(t_{k})=\mathbf{P}_{\text{high}}\bm{\epsilon}(t_{k})$ be the high-frequency component of the error at time $t_{k}$. Applying the projection operator $\mathbf{P}_{\text{high}}$ to both sides of the error evolution equation (Eq. (A.20)):
$\mathbf{P}_{\text{high}}\bm{\epsilon}(t_{k+1})$ $\approx\mathbf{P}_{\text{high}}(\mathbf{A}_{k}\bm{\epsilon}(t_{k})+\bm{e}_{\text{model}}(t_{k}))$ (A.23)
$\bm{\epsilon}_{\text{high}}(t_{k+1})$ $\approx\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})+\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k}).$ (A.24)
Taking the norm and applying the triangle inequality:
$\|\bm{\epsilon}_{\text{high}}(t_{k+1})\|\leq\|\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})\|+\|\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k})\|.$ (A.25)
We make the simplifying approximation that the amplification of high-frequency error is primarily driven by existing high-frequency error components, consistent with the assumption in Eq. (A.21). That is, we consider the first term on the right-hand side to be bounded in relation to $\|\bm{\epsilon}_{\text{high}}(t_{k})\|$:
$\|\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})\|\leq C_{A}\gamma_{\text{high}}\|\bm{\epsilon}_{\text{high}}(t_{k})\|.$ (A.26)
This step uses the assumption Eq. (A.21) by focusing on the impact of $\mathbf{A}_{k}$ on the high-frequency part of $\bm{\epsilon}(t_{k})$, or assumes that the cross-term effects (low-frequency error contributing to high-frequency error via $\mathbf{A}_{k}$) are implicitly captured or subdominant compared to the self-amplification of high-frequency error. By the theorem’s assumption, the single-step high-frequency model error is bounded:
$\|\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k})\|\leq\delta_{\text{HF}}.$ (A.27)
Substituting these bounds into Eq. (A.25), we obtain the recursive inequality for the norm of the high-frequency error:
$\|\bm{\epsilon}_{\text{high}}(t_{k+1})\|\leq C_{A}\gamma_{\text{high}}\|\bm{\epsilon}_{\text{high}}(t_{k})\|+\delta_{\text{HF}}.$ (A.28)
Let $G=C_{A}\gamma_{\text{high}}$ and $x_{k}=\|\bm{\epsilon}_{\text{high}}(t_{k})\|$. The inequality becomes:
$x_{k+1}\leq Gx_{k}+\delta_{\text{HF}}.$ (A.29)
Unrolling this recursion for $M$ steps:
$x_{1}$ $\leq Gx_{0}+\delta_{\text{HF}}$
$x_{2}$ $\leq Gx_{1}+\delta_{\text{HF}}\leq G(Gx_{0}+\delta_{\text{HF}})+\delta_{\text{HF}}=G^{2}x_{0}+(G+1)\delta_{\text{HF}}$
$\vdots$
$x_{M}$ $\leq G^{M}x_{0}+(G^{M-1}+G^{M-2}+\dots+G^{1}+G^{0})\delta_{\text{HF}}$
$x_{M}$ $\leq G^{M}x_{0}+\delta_{\text{HF}}\sum_{i=0}^{M-1}G^{i}.$
The sum is a geometric series. Given the condition $G=C_{A}\gamma_{\text{high}}\neq 1$:
$\sum_{i=0}^{M-1}G^{i}=\frac{G^{M}-1}{G-1}.$ (A.30)
Substituting this back, and replacing $x_{M}$ with $\|\bm{\epsilon}_{\text{high}}(t_{M})\|$, $x_{0}$ with $\|\bm{\epsilon}_{\text{high}}(t_{0})\|$, and $G$ with $C_{A}\gamma_{\text{high}}$, we arrive at the bound:
$\|\bm{\epsilon}_{\text{high}}(t_{M})\|\leq(C_{A}\gamma_{\text{high}})^{M}\|\bm{\epsilon}_{\text{high}}(t_{0})\|+\delta_{\text{HF}}\frac{(C_{A}\gamma_{\text{high}})^{M}-1}{C_{A}\gamma_{\text{high}}-1}.$ (A.31)
This completes the proof of Theorem A.1. ∎
A.3 Discussion and Association with Turb-L1
Theorem A.1 (Eq. (A.22)) highlights two primary sources of high-frequency error accumulation in long-term autoregressive predictions:
  1. 1.
    Amplification of initial high-frequency error: The term $(C_{A}\gamma_{\text{high}})^{M}\|\bm{\epsilon}_{\text{high}}(t_{0})\|$ shows that any initial high-frequency error $\|\bm{\epsilon}_{\text{high}}(t_{0})\|$ is amplified by the system’s dynamics, characterized by $C_{A}\gamma_{\text{high}}$. If $C_{A}\gamma_{\text{high}}>1$, this term can grow exponentially with the number of prediction steps $M$.
  2. 2.
    Accumulation of single-step model high-frequency error: The term $\delta_{\text{HF}}\frac{(C_{A}\gamma_{\text{high}})^{M}-1}{C_{A}\gamma_{\text{high}}-1}$ shows that the single-step high-frequency error $\delta_{\text{HF}}$ introduced by the model at each step also accumulates and is amplified by the system dynamics. This term can also grow exponentially if $C_{A}\gamma_{\text{high}}>1$.
For systems like turbulence, which are inherently multi-scale and sensitive to high-frequency details (often exhibiting $C_{A}\gamma_{\text{high}}>1$ due to phenomena like energy cascade and vortex stretching), controlling the growth of high-frequency error is crucial for long-term stability and fidelity. The theorem implies that minimizing the single-step high-frequency model error, $\delta_{\text{HF}}$, is critical.
The Turb-L1 architecture is specifically designed to address this challenge. Traditional deep learning models often suffer from Spectral Bias, leading to poor representation and prediction of high-frequency components, thereby resulting in a relatively large $\delta_{\text{HF}}$. The Hierarchical Dynamics Synthesis (HDS) mechanism within Turb-L1, operating in a multi-grid framework, aims to:
  • Explicitly capture and process high-frequency information using operators sensitive to local, high-frequency variations (e.g., convolutions with small receptive fields).
  • Synthesize dynamics across different frequency bands (high, middle, and low) and then aggregate them, ensuring that high-frequency details are preserved and correctly interact with larger-scale dynamics.
  • The multi-grid architecture facilitates learning across different scales, which is beneficial for representing the full spectrum of turbulent flows.
By effectively mitigating spectral bias, Turb-L1 strives to achieve a significantly smaller $\delta_{\text{HF}}$. According to Eq. (A.22), a smaller $\delta_{\text{HF}}$ directly leads to a slower accumulation of high-frequency error, enhancing the long-term stability and fidelity of turbulence predictions. This theoretical underpinning explains the superior performance of Turb-L1 observed in experiments, particularly in accurately reproducing the enstrophy spectrum at high wavenumbers and maintaining structural details over extended prediction horizons.
Appendix B Dataset Details
In this work, we evaluate our proposed method, Turb-L1, on two canonical benchmark datasets for two-dimensional (2D) isotropic turbulence. These datasets are specifically chosen to assess the model’s capability to capture complex, chaotic fluid dynamics under different physical regimes: one with external forcing and another exhibiting natural decay. All simulations are discretized on a $128\times 128$ uniform spatial grid.
B.1 Dataset I: Forced Isotropic Turbulence
This dataset is based on the Navier-Stokes equation benchmark introduced by Li et al. [20] in their work on Fourier Neural Operators (FNO). It simulates the evolution of vorticity in a 2D viscous, incompressible fluid on the unit torus $(0,1)^{2}$. The governing equations in vorticity form are:
$\frac{\partial w(\mathbf{x},t)}{\partial t}+\mathbf{u}(\mathbf{x},t)\cdot\nabla w(\mathbf{x},t)$ $=\nu\Delta w(\mathbf{x},t)+f(\mathbf{x})\quad\forall\mathbf{x}\in(0,1)^{2},t\in(0,T]$ (A.32)
$\nabla\cdot\mathbf{u}(\mathbf{x},t)$ $=0\quad\forall\mathbf{x}\in(0,1)^{2},t\in[0,T]$ (A.33)
$w(\mathbf{x},0)$ $=w_{0}(\mathbf{x})\quad\forall\mathbf{x}\in(0,1)^{2}$ (A.34)
where $w$ is the vorticity field ($w=\nabla\times\mathbf{u}$), $\mathbf{u}$ is the velocity field, $w_{0}$ is the initial vorticity field, $\nu$ is the kinematic viscosity coefficient, and $f(\mathbf{x})$ is a time-independent external forcing function.
In our study, we adopt the setup from Li et al. [20] and specifically focus on the case where the kinematic viscosity is set to $\nu=1\times 10^{-5}$. The initial vorticity field, $w_{0}$, is generated by sampling from a zero-mean Gaussian random field whose covariance operator is $(-\Delta+\tau^{2}I)^{-\alpha/2}$. Here, $\Delta$ denotes the Laplacian operator, $I$ is the identity operator, and the parameters $\tau$ and $\alpha$ control the spectral shape, such that the initial energy spectrum density $E(k)$ follows $E(k)\sim(k^{2}+\tau^{2})^{-\alpha}$. The external forcing term $f(\mathbf{x})$ is a fixed force, predominantly concentrated in the low-wavenumber regime, designed to continuously inject energy into the system. No additional drag effects are considered in the simulations.
This dataset is employed to evaluate the model’s long-term prediction capabilities and its ability to capture chaotic dynamics in the presence of sustained external energy input. In the experiments presented in this paper, we extend the prediction horizon for this dataset up to 19 time steps.
B.2 Dataset II: Decaying Isotropic Turbulence (McWilliams, 1984)
This dataset simulates the classic phenomenon of 2D decaying isotropic turbulence. The setup for the initial conditions follows the seminal work by McWilliams [24], titled "The emergence of isolated coherent vortices in turbulent flow." Unlike Dataset I, in this scenario, the fluid system evolves without any external forcing after an initial perturbation; its energy naturally decays over time.
The initial flow field is characterized by the power spectrum of its streamfunction $\psi(\mathbf{x})$. Specifically, the squared magnitude of the Fourier coefficients of the initial streamfunction, $|\hat{\psi}(\mathbf{k})|^{2}$ (where $\mathbf{k}$ is the wavevector), adheres to the following spectral distribution:
$|\hat{\psi}(\mathbf{k})|^{2}\sim k^{-1}(\tau_{0}^{2}+(k/k_{0})^{4})^{-1}$ (A.35)
where $k=|\mathbf{k}|$ is the wavenumber magnitude, $k_{0}$ is a characteristic wavenumber, and $\tau_{0}$ is a parameter that shapes the spectrum. These initial conditions are designed to facilitate a slow energy decay and allow the system to self-organize into coherent vortex structures during its evolution. The enstrophy density spectrum also evolves to exhibit features akin to a Kolmogorov energy cascade.
This dataset is a crucial benchmark for evaluating a model’s ability to capture long-term dynamical behavior, energy dissipation mechanisms, and the formation and evolution of complex structures (such as isolated coherent vortices) in a purely self-evolving system. In our experiments, we extend the prediction horizon for this dataset up to 99 time steps to thoroughly assess the long-term stability and fidelity of the models.
Appendix C Metrics
C.1 $L^{2}$ Error
Measures the $L^{2}$ norm of the error field, which is the difference between the predicted field ($\hat{\omega}(\cdot,t)$) and the ground truth field ($\omega(\cdot,t)$).
$L^{2}\text{ Error}(t)=\|\hat{\omega}(\cdot,t)-\omega(\cdot,t)\|_{L^{2}}=\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\hat{\omega}_{i,j}(t)-\omega_{i,j}(t))^{2}}$ (A.36)
C.2 Structural Similarity Index Measure (SSIM)
Assesses the perceptual similarity between two fields (prediction $u$, ground truth $v$) considering local means ($\mu$), variances ($\sigma^{2}$), and covariance ($\sigma_{uv}$).
$\text{SSIM}(u,v)=\frac{(2\mu_{u}\mu_{v}+c_{1})(2\sigma_{uv}+c_{2})}{(\mu_{u}^{2}+\mu_{v}^{2}+c_{1})(\sigma_{u}^{2}+\sigma_{v}^{2}+c_{2})}$ (A.37)
C.3 Relative $L^{2}$ Error
Normalizes the $L^{2}$ norm of the error field by the $L^{2}$ norm of the ground truth field.
Relative $\text{Relative }L^{2}\text{ Error}(t)=\frac{\|\hat{\omega}(\cdot,t)-\omega(\cdot,t)\|_{L^{2}}}{\|\omega(\cdot,t)\|_{L^{2}}}=\frac{\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\hat{\omega}_{i,j}(t)-\omega_{i,j}(t))^{2}}}{\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\omega_{i,j}(t))^{2}}}$ (A.38)
C.4 Enstrophy Spectrum (Density), $E_{Z}(k,t)$
[…]
$E_{Z}(k,t)\approx\frac{1}{\Delta k}\sum_{k\leq|\mathbf{k}^{\prime}|<k+\Delta k}\frac{1}{2A}|\hat{\omega}(\mathbf{k}^{\prime},t)|^{2}\cdot(2\pi k^{\prime})$ (A.39)
C.5 Normalized Spectral Error
Quantifies the relative difference between the predicted enstrophy spectrum ($E_{Z,\text{pred}}(k,t)$) and the ground truth enstrophy spectrum ($E_{Z,\text{GT}}(k,t)$) at each wavenumber $k$.
Normalized Spectral $\text{Normalized Spectral Error}(k,t)=\frac{E_{Z,\text{pred}}(k,t)-E_{Z,\text{GT}}(k,t)}{E_{Z,\text{GT}}(k,t)}$ (A.40)
Appendix D Experimental Configurations
This section details the model hyperparameters, training procedures, data handling, and logging configurations for the experiments conducted on both datasets.
D.1 Configurations for Dataset I: Forced Isotropic Turbulence
D.1.1 Model Architectural Hyperparameters (Dataset I)
Table LABEL:tab:model_hyperparams_d1 summarizes the key architectural hyperparameters. For models with extensive parameter sets (e.g., CNO), please refer to the original publications or supplementary code.
Table 2: Architectural Hyperparameters for Dataset I (Forced Isotropic Turbulence).
Model Name Parameter Value(s)
Turb-L1 Input Shape (T,C,H,W) [1, 1, 128, 128]
Spatial Hidden Dim 128
Output Channels 1
Temporal Hidden Dim 256
Num Spatial Layers 4
Num Temporal Layers 8
FNO Modes (modes1, modes2) 8, 8
Width 64
Input Channels (C_in) 1
Output Channels (C_out) 1
SimVP Input Shape (T,C,H,W) [1, 1, 64, 64]
Hidden Dim S (hid_S) 128
Hidden Dim T (hid_T) 256
Output Dim 1
CNO Input Channels (in_dim) 1
Input Size (H, W) 64, 64
Num Layers (N_layers) 4
Channel Multiplier 32
Latent Lift Proj Dim 64
Activation ’cno_lrelu’
MGNO Num Layers 6
Output/Input Channels 1
Height / Width 128 / 128
LSM Model Dim (d_model) 64
Num Tokens / Basis 4 / 16
Patch Size ’4,4’
U-Net Input/Output Channels 1 / 1
Kernel Size 3
Dropout Rate 0.5
ResNet Input/Output Channels 1 / 1
Dropout Rate 0.5
ViT Image Size / Patch Size 128 / 4
Input/Output Channels 1 / 1
Embedding Dim / Depth 256 / 1
PastNet Input Shape (T,C,H,W) [1, 1, 64, 64]
Hidden Dim T (hid_T) 256
Inception Kernels [3, 5, 7, 11]
ConvLSTM Input Dim / Hidden Dims 1 / [64, 64]
Num Layers 2
KNO Input/Output Channels 1 / 1
Operator Size (op_size) 64
Modes X / Y 10 / 10
D.1.2 Training Configurations (Dataset I)
Table 3 details the training configurations. All models were trained for 500 epochs with a seed of 42.
Table 3: Training Configurations for Dataset I.
Model Name Batch Size Learning Rate Parallel Method
Turb-L1 20 0.001 DistributedDataParallel
FNO 1 0.001 DataParallel
SimVP 20 0.001 DistributedDataParallel
CNO 20 0.001 DistributedDataParallel
MGNO 20 0.0005 DistributedDataParallel
LSM 20 0.001 DistributedDataParallel
U-Net 20 0.001 DistributedDataParallel
ResNet 20 0.001 DistributedDataParallel
ViT 20 0.001 DistributedDataParallel
PastNet 20 0.001 DistributedDataParallel
ConvLSTM 20 0.001 DistributedDataParallel
KNO 500 0.001 DataParallel
D.2 Configurations for Dataset II: Decaying Isotropic Turbulence
D.2.1 Model Architectural Hyperparameters (Dataset II)
Table LABEL:tab:model_hyperparams_d2 summarizes key architectural hyperparameters for Dataset II.
Table 4: Architectural Hyperparameters for Dataset II (Decaying Isotropic Turbulence).
Model Name Parameter Value(s)
Turb-L1 Input Shape (T,C,H,W) [1, 1, 128, 128]
Spatial Hidden Dim 128
Output Channels 1
Temporal Hidden Dim 256
Num Spatial Layers 4
Num Temporal Layers 8
FNO Modes (modes1, modes2) 16, 16
Width 64
Input Channels (C_in) 1
Output Channels (C_out) 1
SimVP Input Shape (T,C,H,W) [1, 1, 128, 128]
Hidden Dim S (hid_S) 64
Hidden Dim T (hid_T) 128
Output Dim 1
CNO Input Channels (in_dim) 1
Input Size (H, W) 128, 128
Num Layers (N_layers) 4
Channel Multiplier 32
Latent Lift Proj Dim 64
Activation ’cno_lrelu’
MGNO Num Layers 6
Output/Input Channels 1
Height / Width 128 / 128
LSM Model Dim (d_model) 64
Num Tokens / Basis 4 / 16
Patch Size ’4,4’
U-Net Input/Output Channels 1 / 1
Kernel Size 3
Dropout Rate 0.5
ResNet Input/Output Channels 1 / 1
Dropout Rate 0.5
ViT Image Size / Patch Size 128 / 4
Input/Output Channels 1 / 1
Embedding Dim / Depth 256 / 1
PastNet Input Shape (T,C,H,W) [1, 1, 128, 128]
Hidden Dim T (hid_T) 256
Inception Kernels [3, 5, 7, 11]
ConvLSTM Input Dim / Hidden Dims 1 / [64, 64]
Num Layers 2
KNO Input/Output Channels 1 / 1
Operator Size (op_size) 128
Modes X / Y 10 / 10
D.2.2 Training Configurations (Dataset II)
Training configurations (batch size, learning rate, epochs, seed, parallel method) for Dataset II were identical to those for Dataset I (see Table 3), unless specified otherwise.
Appendix E Ablation Study
E.1 Ablation Study Setup for Decaying Isotropic Turbulence

### v3

@@ 5.2 Main results
Table I presents the performance achieved across different baselines over various prediction steps. We summarize the key takeaways as follows:
Takeaway ➊: Turb-L1 consistently achieves state-of-the-art performance across diverse benchmarks and prediction steps.  Across both Dataset I and II, Turb-L1 achieves significantly lower $L^{2}$ errors compared to all three mainstream method categories (,   and Y). Notably, on Dataset I, Turb-L1 demonstrates exceptional performance with an error of only 0.6171, substantially outperforming the second-best baseline NMO (2.1923). This superior performance remains consistently observable in Dataset II (0.1257 vs. 0.1873). More generally, Turb-L1 attains overall error reductions ranging from $43.66\%\sim 92.29\%$ across different prediction steps, conclusively validating our Turb-L1 effectiveness for long-term forecasting.
Takeaway ➋: Turb-L1 demonstrates excellent visual fidelity and physical consistency in long-term predictions. It effectively overcomes over-smoothing and simulation collapse, issues that spectral bias causes. As Figure 3 illustrates, Turb-L1’s predictions closely match the ground truth even after 99 time steps, clearly reproducing complex vortex structures and fine flow features. This performance stems from its core Hierarchical Dynamics Synthesis mechanism, which precisely captures full-spectrum dynamic information, particularly high-frequency details. In contrast, FNO exhibits significant over-smoothing due to spectral bias and loses critical high-frequency information. SimVP, on the other hand, experiences physics collapse from accumulated high-frequency error amplification, producing unrealistic artifacts. Turb-L1’s SSIM and relative L2 error curves further confirm its significant advantages in maintaining structural similarity and long-term prediction accuracy. This fully demonstrates the method’s robustness and high fidelity in tracking complex multi-scale turbulence dynamics.
@@ 5.4 Turb-L1’s Superiority in Enstrophy Spectrum Prediction
Figure 8: Ablation study on the long-term predictive performance of Turb-L1 and its variants. The plot shows the Mean Relative Error accumulated over 100 autoregressive prediction steps. The full Turb-L1 model (solid blue line) significantly outperforms all ablated versions, maintaining a very low and stable error growth rate. Removing the Multi-Grid architecture (dashed green line, w/o MG) leads to a noticeable increase in error, though it remains more stable than other variants. Crucially, removing the core HDS mechanism (dot-dashed orange line, w/o HDS) or relying on only a single frequency band (Low Freq Only or High Freq Only) results in catastrophic performance degradation, with errors rapidly accumulating. These results demonstrate that the synergy between the multi-grid architecture and the Hierarchical Dynamics Synthesis (HDS) mechanism is essential for achieving the long-term stability and accuracy of Turb-L1.
TABLE II: Ablation study on the McWilliams 2D decaying turbulence dataset. We evaluate the performance of Turb-L1 against its variants by removing key components: the Hierarchical Dynamics Synthesis (HDS) module and the Multi-Grid (MG) architecture. The variants ‘w/o High-Freq.‘ and ‘w/o Low-Freq.‘ represent models trained without the respective frequency-processing paths. Lower Relative L2 Error and higher SSIM indicate better performance. The results clearly demonstrate that every component is crucial for achieving state-of-the-art performance.
Model Prediction Steps
Step 50 Step 99
R. $L^{2}(\downarrow)$ SSIM $(\uparrow)$ R. $L^{2}(\downarrow)$ SSIM $(\uparrow)$
Turb-L1 (w/o HDS) 1.2240 0.0451 1.4032 0.0378
Turb-L1 (w/o MG) 0.3001 0.8475 0.5825 0.5596
w/o High-Freq. 1.0494 0.1098 1.1620 0.0519
w/o Low-Freq. 1.1116 0.0822 1.2493 0.0396
Turb-L1 (Full) 0.1622 0.9496 0.3772 0.7797
@@
5.5 Ablation Study
5.5.1 Ablation Study Setup for Decaying Isotropic Turbulence